# ABD: Default–Exception Abduction in Finite First-Order Worlds

### v1

@@ G.3 Beats-Gold Breakdown
T1 0/30  (0.0%) 1/28  (3.6%) 9/143  (6.3%)
T2 31/318  (9.7%) 4/422  (0.9%) 13/123 (10.6%)
T3 22/225  (9.8%) 4/310  (1.3%) 26/126 (20.6%)
T4 23/309  (7.4%) 5/401  (1.2%) 32/140 (22.9%)
T5 32/175 (18.3%) 8/185  (4.3%) 23/141 (16.3%)
T6 48/83  (57.8%)
T7 5/138  (3.6%)
T6 in ABD-Skeptical stands out at 57.8%: the planted gold for TH3 ($P(x)\land\lnot\mathsf{Ab}(x)\to\exists y\,R(x,y)$) yields a generous gold cost under worst-case completion, and models often find tighter formulas. ABD-Partial has universally low beats-gold rates ($\leq 4.3\%$), reflecting good gold-formula design for existential completion.
@@
$[0,5)$ 0/6  (0.0%)
$[5,10)$ 24/161 (14.9%)
$[10,15)$ 19/132 (14.4%) 9/370  (2.4%) 52/397 (13.1%)
$[15,20)$ 51/397 (12.8%) 6/438  (1.4%) 33/123 (26.8%)
$[20,30)$ 10/406  (2.5%) 5/420  (1.2%) 44/183 (24.0%)
$[30,+)$ 28/122 (23.0%) 1/113  (0.9%) 3/24  (12.5%)
In ABD-Full, the highest beats-gold rate occurs for the most complex golds (AST $\geq 30$, 23.0%), where the planted formula contains redundancy that models can trim. In ABD-Skeptical, mid-range golds (AST 15–30) are beaten most often (24–27%), suggesting these formulas are complex enough to leave room for simpler alternatives under worst-case semantics.
@@
The largest single-instance margin is +14+14 in ABD-Skeptical (instance abd_skeptical_v2_TH3_00, T6): five models independently discover a formula with train cost 7 vs. gold cost 21. In ABD-Full, the largest margin is +7+7 (GPT-5.2 on ABD_FULL_TH7_021, T2: predicted cost 12 vs. gold cost 19). ABD-Partial margins are smaller (max +4+4), consistent with the tight gold design under existential completion.
@@ H.1 Worked Examples per Regime
Theory T2: $\exists y(R(x,y)\land P(y))\land\lnot\mathsf{Ab}(x)\to\exists z(S(x,z)\land Q(z))$. 10 training worlds, 5 holdout worlds. 5 models produce robust formulas; 2 are brittle.
Robust (Grok4.1f, AST = 37):
(exists y (exists z (and (not (= y z))
  (R x y) (R x z) (or (P y) (P z))
  (forall w (or (not (R x w)) (or (= w y) (= w z)))))))
Train cost = 13 (gap 0.10/world); holdout cost = 9, all valid.
Interpretation: $x$ has exactly two $R$-successors, at least one of which satisfies $P$. The cardinality constraint generalises because the domain structure is stable.
@@
Same theory as above, but under partial observation. 8 training worlds, 5 holdout worlds. 1 robust, 6 brittle predictions.
@@
Theory T1: $\exists y(R(x,y)\land P(y))\land\lnot\mathsf{Ab}(x)\to Q(x)$. 5 training worlds, 5 holdout worlds. 3 robust, 3 brittle predictions.
@@
Train cost = 5; holdout: catastrophic failure.
Failure: the uniqueness constraint ($\forall z$-clause forces $z=y$) insists that $x$ has exactly one $R$-related $P$-element. Holdout worlds readily contain elements with multiple $P$-witnesses, invalidating the formula under skeptical completion.
@@ H.2 Common Failure Patterns
Table H.1: Failure Mode Breakdown. Aggregated across all abduction tasks. Parse = formula extraction failed; Invalid = parsed but not valid on all train worlds; Valid = train-valid predictions. Best values bold.
Model Valid% Invalid% Parse% Missing%
Opus-4.6 98.8% 0.5% 0.0% 0.7%
Grok4 89.3% 10.7% 0.0% 0.0%
GPT-5.2 92.3% 7.7% 0.0% 0.0%
Gemini-3.1 98.0% 2.0% 0.0% 0.0%
Gemini-3 76.8% 23.2% 0.0% 0.0%
Grok4.1f 94.8% 5.2% 0.0% 0.0%
DSR 94.7% 5.3% 0.0% 0.0%
Kimi-K2t 69.5% 30.5% 0.0% 0.0%
Hermes4 2.3% 97.7% 0.0% 0.0%
GPT-4o 19.8% 80.2% 0.0% 0.0%
@@
    Parse error: formula does not parse as a valid S-expression.
@@
Table H.2: Failure-mode counts across all 5,971 model predictions.
@@
Parse error 0 4 0
All invalid (train) 332 406 315
Partial invalid (train) 241 241 41
Brittle (holdout) 254 452 465
of which catastrophic 73 99 265
Parsimony inflation 306 189 72
@@
Parse errors are negligible (4 total, all from ABD-Partial). The dominant failure mode in ABD-Partial is all invalid on train (406), reflecting the difficulty of satisfying existential completion across 8+ worlds. Brittle failures increase monotonically with scenario difficulty: ABD-Skeptical has the most brittle predictions (465) and the highest catastrophic rate (265/465 = 57% of brittle predictions fail on $\geq$50% of holdout worlds, vs. 29% for ABD-Full). Parsimony inflation moves in the opposite direction (306 in Full, 72 in Skeptical): models in Skeptical tend to fail validity outright rather than produce valid-but-expensive formulas.

### v2

@@ G.3 Beats-Gold Breakdown
T1 4/42 (9.5%) 5/42 (11.9%) 14/239 (5.9%)
T2 63/437 (14.4%) 32/604 (5.3%) 36/186 (19.4%)
T3 62/323 (19.2%) 16/431 (3.7%) 53/185 (28.6%)
T4 77/417 (18.5%) 48/573 (8.4%) 75/219 (34.2%)
T5 63/247 (25.5%) 30/258 (11.6%) 58/208 (27.9%)
T6 86/129 (66.7%)
T7 9/230 (3.9%)
T6 in ABD-Skeptical stands out at 66.7%: the planted gold for TH3 ($P(x)\land\lnot\mathsf{Ab}(x)\to\exists y\,R(x,y)$) yields a generous gold cost under worst-case completion, and models often find tighter formulas. Across ABD-Full and ABD-Skeptical, beats-gold is now common for several theories, especially T3–T6. ABD-Partial remains the least permissive regime, but once GPT-5.4 is included some theories still reach 8–12%, so the gap between scenarios is quantitative rather than absolute.
@@
$[0,5)$ 0/9 (0.0%)
$[5,10)$ 51/254 (20.1%)
$[10,15)$ 45/198 (22.7%) 33/510 (6.5%) 111/629 (17.6%)
$[15,20)$ 112/546 (20.5%) 60/636 (9.4%) 67/187 (35.8%)
$[20,30)$ 59/565 (10.4%) 30/595 (5.0%) 91/279 (32.6%)
$[30,+)$ 53/157 (33.8%) 5/155 (3.2%) 11/38 (28.9%)
In ABD-Full, the highest beats-gold rate occurs for the most complex golds (AST $\geq 30$, 34.1%), where the planted formula contains redundancy that models can trim. In ABD-Skeptical, beats-gold is strongest for mid-range and moderately large golds, peaking at AST 15–20 (35.8%) and remaining high at AST 20–30 (32.6%), suggesting these formulas are complex enough to leave room for simpler alternatives under worst-case semantics. ABD-Partial remains comparatively low across AST bins, never exceeding 9.2%.
@@
The largest single-instance margin is +14+14 in ABD-Skeptical (instance abd_skeptical_v2_TH3_00, T6): nine models independently discover a formula with train cost 7 vs. gold cost 21. In ABD-Full, the largest margin is now +8+8 (GPT-5.4 on ABD_FULL_TH2_094, T1: predicted cost 12 vs. gold cost 20). ABD-Partial reaches the same maximum +8+8 (GPT-5.4 on ABD_PARTIAL_TH7_024b, T2: predicted cost 10 vs. gold cost 18), showing that once GPT-5.4 is included, substantial single-instance improvements over gold are no longer confined to the other two regimes.
@@ H.1 Worked Examples per Regime
Theory T2: $\exists y(R(x,y)\land P(y))\land\lnot\mathsf{Ab}(x)\to\exists z(S(x,z)\land Q(z))$. 10 training worlds, 5 holdout worlds. 2 models produce robust formulas; 7 are brittle.
Robust (Opus-4.6, AST = 8):
(exists y (and (R x y) (P y)))
Train cost = 34 (gap 2.20/world); holdout cost = 25, all valid.
Interpretation: the formula marks any $x$ with an $R$-related $P$-witness as abnormal. It is less cost-efficient on the training set than specialized alternatives, but it generalizes because it avoids accidental cardinality assumptions about the local $R$-neighborhood.
@@
Same theory as above, but under partial observation. 8 training worlds, 5 holdout worlds. 1 robust, 8 brittle predictions.
@@
Theory T1: $\exists y(R(x,y)\land P(y))\land\lnot\mathsf{Ab}(x)\to Q(x)$. 5 training worlds, 5 holdout worlds. 6 robust, 3 brittle predictions.
@@
Train cost = 5; holdout: 1/5 worlds invalid (80% valid).
Failure: the uniqueness constraint ($\forall z$-clause forces $z=y$) insists that $x$ has exactly one $R$-related $P$-element. Under skeptical completion, even a single holdout world with multiple $P$-witnesses is enough to invalidate the rule.
@@ H.2 Common Failure Patterns
Table H.1: Failure Mode Breakdown. Aggregated across all abduction tasks. Parse = unrepaired parse failure; Repair = trailing right-parenthesis repair applied before evaluation; Invalid = parsed but not valid on all train worlds; Valid = train-valid predictions. Best values bold.
Model Valid% Invalid% Parse% Repair% Missing%
Opus-4.6 98.8% 0.5% 0.0% 0.0% 0.7%
GPT-5.2 92.2% 6.0% 1.8% 0.2% 0.0%
Gemini-3.1 98.0% 2.0% 0.0% 0.0% 0.0%
Grok4.1f 95.2% 3.9% 0.9% 0.7% 0.0%
DSR 95.9% 3.9% 0.2% 1.6% 0.0%
GPT-5.4 85.2% 14.0% 0.8% 6.7% 0.0%
Grok4 89.4% 6.2% 4.4% 0.2% 0.0%
Gemini-3 76.9% 21.7% 1.4% 0.0% 0.0%
Kimi-K2t 71.5% 12.2% 16.3% 6.0% 0.0%
Hermes4 2.3% 89.8% 8.0% 0.0% 0.0%
GPT-4o 19.8% 77.7% 2.5% 0.0% 0.0%
@@
    Auto-repaired parse: the only parse failure is missing trailing right parentheses, so we conservatively close the suffix and evaluate the repaired formula.
@@
    Parse error: formula still does not parse as a valid S-expression after the conservative suffix-repair attempt.
  3. 3.
@@
  7. 7.
[…]
Table H.2: Failure-mode counts across all 6,367 model predictions.
@@
Auto-repaired parse 47 26 18
Parse error 87 43 86
All invalid (train) 208 343 218
Partial invalid (train) 288 270 54
Brittle (holdout) 357 583 571
of which catastrophic 121 174 332
Parsimony inflation 292 183 73
@@
Auto-repaired parses are not rare: there are 93 across 6,571 predictions, compared with 217 unrepaired parse errors. Syntax failures therefore remain a minority, but they are not negligible. The dominant failure mode in both ABD-Partial and ABD-Skeptical is brittleness on holdout (603 and 571 cases, respectively), rather than train-time invalidity. Catastrophic brittleness is concentrated in ABD-Skeptical (332/571 = 58% of brittle predictions fail on at least half of holdout worlds). Parsimony inflation moves in the opposite direction, peaking in ABD-Full (305 cases) and becoming much rarer in ABD-Skeptical (73): under universal completion, models more often fail validity outright than remain valid with inflated cost.